# Pay-per-search models are abstention models

### v1

@@ 5.3 Analysis III: Out-of-Distribution Performance
Method Natural Questions TriviaQA
[…]
otc 2.09 54.35 99.04 8.32 4.07 71.43 96.95 7.11
MASH w/ otc-st $18.25$ 51.24 79.95 51.63 30.52 67.61 77.53 51.55
DPO 24.4 - 77.38 68.23 41.6 - 71.71 66.23
Results: Table 5 reports our results (NaturalQA and 2Wiki models are in Appendix D). MASH generalizes better than the OTC (higher Accuracy and Delta values), which abstains on nearly all questions out-of-distribution. MASH also reports better Abs(0) performance that DPO but lower Delta. We attribute this to MASH generalizing more conservatively out-of-domain. With 2Wiki, which exclusively contains two-hop questions, MASH generalizes relatively well to HotPotQA but fails on single-hop datasets. We argue that, under poor out-of-distribution accuracy generalization, abstention and invoking search tools is the more ideal decision. With search enabled, our HotPotQA-trained MASH model attains $26.43\%$ higher accuracy than DPO, which is limited to abstention.
@@ 8 Acknowledgments
Appendix A Search Tool Use
@@ A.2 Inference Algorithm
Question $q$, language model $\pi_{\theta}$, retriever $H$
Maximum search budget $L$
Trajectory $\tau$
@@ A.3 Warm-Start
Evaluation of trajectories is done with an LLM judge, in this case Qwen2.5-72B-Instruct (Qwen et al., 2025). We follow the same procedure we use to evaluate abstention model outputs, described in Section B.1. If a trajectory is deemed correct, we swap its generated answer with the ground-truth answer for the target dataset to align answers with the dataset format, as we use exact match as the reward.
@@
A.4 Datasets
We run training experiments on three knowledge-intensive datasets – the single-hop dataset Natural Questions (Kwiatkowski et al., 2019), and multi-hop datasets HotPotQA (Yang et al., 2018) and 2WikiMultiHopQA (Ho et al., 2020). We additionally use the single-hop TriviaQA dataset as part of our out-of-distribution evaluations. For Natural Questions, we use the official splits for training, validation and test. For HotPotQA, 2WikiMultiHopQA and TriviaQA, the official test splits do not contain answers. As a result, we use their official development/validation sets for the purpose of test and construct our own validation sets by sub-sampling from the training set with a 90/10 split.
Additionally, as noted in the main text, we filter out the “comparison” and “bridge-comparison” questions from 2WikiMultiHopQA, as these questions are each binary choice questions with heavily skewed answer distributions, causing models to exploit dataset distributions in practice.
Input Prompt: Answer the given question. You should first have a reasoning process in mind and then provides the answer. Show your reasoning in <<think>> <</think>> tags and return the final answer in <<answer>> <</answer>> tags, for example <<answer>> Beijing <</answer>>. Question: <<question>>
@@ A.4 Datasets
Input Prompt: Answer the given question. You must conduct reasoning between <<think>> and <</think>> every time you get new information. After reasoning, if you find you lack some knowledge, you can call a search engine by <<search>> query <</search>> and it will return the top searched results between <<document>> and <</document>>. You need to make every search call count and gain helpful results. If you find no further external knowledge is needed, you can directly provide the answer inside <<answer>> and <</answer>>, without detailed illustrations. For example, <<answer>> Beijing <</answer>>. Question: <<question>>
@@
Input Prompt: Answer the given question. You must conduct reasoning between <<think>> and <</think>> every time you get new information. After reasoning, if you find you lack some knowledge, you can ask a question to a search engine by <<search>> query <</search>> and it will return the top searched results between <<document>> and <</document>>. A search query should be an atomic question asking about one, single piece of information. Example 1: Question: “Who was born first, Clint Eastwood or Harrison Ford?” Valid Queries: “<<search>>Clint Eastwood birth date<</search>>” and “<<search>>Harrison Ford birth date<</search>>”. The query “<<search>>Clint Eastwood and Harrison Ford birth date<</search>>” is invalid. The query “ <<search>> Clint Eastwood birth date Harrison Ford birth date <</search>> is also invalid. Do not pack in multiple questions into one query. Each query should be completely independent. Example 2: Question: “Which is a genus of palms, Zinnia or Butia?” Valid Queries: “<<search>>Zinnia genus classification<</search>>” and “<<search>>Butia genus classification<</search>>”. Example 3: Question: “When did the country where Piltene is located become part of the USSR?” Initial Query: “<<search>>Piltene location<</search>> In each of these examples, you should conduct a search only if you lack the relevant information. Remember, you should decompose questions in your search queries and conduct searches for each atomic question separately. You need to make every search call count and gain helpful results. If you find no further external knowledge is needed, you can directly provide the answer inside <<answer>> and <</answer>>, without detailed illustrations. For example, <<answer>> Beijing <</answer>>. Question: <<question>>
@@
Appendix B Abstention Experiment Details
@@
B.1 Question Accuracy Estimation
We follow the pipeline used by  Yang et al. (2024) to estimate the average accuracies. For a given question $q$, we sample $10$ responses $\{\hat{y_{i}}\}^{10}_{i=1}$ from the untrained model. As all of our experiments are conducted with base models, we perform few-shot prompting. Specifically, for each dataset, we collect correct responses sampled from DeepSeek-V3.1 to $5$ questions sampled from the training set and use these as our few-shot examples. For this component, we perform inference with DeepSeek-V3.1 using a temperature of $1$ and top-$p$ of $0.8$. We likewise perform sampling with Qwen2.5-3B with a temperature of $1$, top-$p$ of $0.8$ and top-$k$ of $50$ to ensure that the base model samples strong outputs and gives a good estimate of knowledge boundaries.
To assess the correctness of a given answer $\hat{y}_{i}$, we first extract a shortform response and then evaluate the accuracy of this extracted response with an LLM judge. We use DeepSeek-V3.1 in both cases using the few-shot prompts of Yang et al. (2024) (shown in Figures 6 and 7), using greedy decoding for replicability.
B.2 Trained Abstention Model Details
@@
B.3 Few-Shot Abstention Prompting Details
@@
B.4 Evaluating Abstention Models

### v2

@@ 4.6 Out-of-Distribution Performance
Results: Table 5 reports our results (NaturalQA and 2Wiki models are in Appendix G). MASH generalizes better than the OTC (higher Accuracy and Delta values), which abstains on nearly all questions out-of-distribution. MASH also reports better Abs(0) performance than DPO, but lower Delta. We attribute this to MASH generalizing more conservatively out-of-domain. With 2Wiki, which only contains multi-hop questions, MASH generalizes well to HotPotQA but performs poorer on single-hop datasets. We argue that, for out-of-distribution generalization, abstention and invoking search tools is the more ideal decision. With search enabled, our HotPotQA-trained MASH model attains $26.43\%$ higher accuracy than DPO, which is limited to abstention.
@@ 7 Acknowledgments
Appendix A Theoretical Analysis
A.1 Preliminaries
In this section, we provide a theoretical analysis of the search behavior the optimal policy for the proxy objective of Equation 1 will display for a given question. Specifically, we will formally demonstrate that the optimal policy must produce a parametric answer to a question $q$ if and only if the maximum achievable expected accuracy when answering $q$ parametrically is greater than or equal to the maximum achievable expected reward when performing searches.
Before starting the analysis proper, we will first make two assumptions:
  1. -
    Assumption 1: We set the KL penalty weight $\beta=0.0$ to simplify the proxy objective and focus solely on the search behavior of an optimal policy maximizing reward.
  2. -
    Assumption 2: We assume that the optimal policy cannot achieve perfect expected accuracy (i.e., average accuracy across multiple samples), for all questions when answering parametrically. In other words, the maximum achievable expected accuracy is bounded for some questions. This follows by definition from our training setup, where an LLM cannot acquire new knowledge beyond the base model’s knowledge boundaries during RL and from prior work, which suggests that standard finetuning approaches on new knowledge risk an increase in hallucinations rather than incorporation of knowledge (Lin et al., 2024; Gekhman et al., 2024).
We then make three definitions to be used in our proofs:
  1. -
    Definition 1: Let $N_{s}(\tau)$ be the number of search calls made by a trajectory $\tau$.
  2. -
    Definition 2: Let $r_{acc}(y,\tau)\in\{0,1\}$ and $r_{help}(q,\tau)$ be defined such that $r_{acc}(y,\tau)\cdot r_{help}(q,\tau)=1$ if $r_{acc}(y,\tau)=1$ and $N_{s}(\tau)=0$. This follows from the definitions set in Section 2.1.
  3. -
    Definition 3: Consider a question-answer pair $(q,y)$. Then, we define:
    $R^{*}_{i}(q,y)=\max_{\theta}E_{\tau\sim\pi_{\theta}(\cdot|q,H)}[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)|N_{s}(\tau)=i]$
    as the maximum achievable expected reward by a policy when performing $i$ searches.
A.2 Proof
Lemma 1:
For a given question-answer pair $(q,y)$, the optimal policy $\pi_{\theta^{*}}$ must answer the question $q$ parametrically (without any searches) if and only if $R^{*}_{0}(q,y)>R^{*}_{i}(q,y)$ for all $i>0$.
Proof of Lemma 1:
We firstly note that the expected reward for any arbitrary policy $\theta$ given a question-answer pair $(q,y)$ can be written as a weighted sum of its expected reward when answering question $q$ with different search counts. We show this below:
$E_{\tau\sim\pi_{\theta}(\cdot|q,H)}$ $[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)]=\Sigma_{\tau}\left[\pi_{\theta}(\tau|q,H)\cdot r_{acc}(y,\tau)\cdot r_{help}(q,\tau)\right]$
$=\Sigma^{\infty}_{i=0}\Sigma_{\tau|N_{s}(\tau)=i}\left[\pi_{\theta}(\tau|q,H)\cdot r_{acc}(y,\tau)\cdot r_{help}(q,\tau)\right]$
$=\Sigma^{\infty}_{i=0}\left[P(N_{s}(\tau)=i|q;\theta)\cdot\Sigma_{\tau|N_{s}(\tau)=i}\left[\frac{\pi_{\theta}(\tau|q,H)}{P(N_{s}(\tau)=i|q;\theta)}\cdot r_{acc}(y,\tau)\cdot r_{help}(q,\tau)\right]\right]$
$=\Sigma^{\infty}_{i=0}\left[P(N_{s}(\tau)=i|q;\theta)\cdot E_{\tau\sim\pi_{\theta}(\cdot|q,H)}\left[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)|N_{s}(\tau)=i\right]\right],$
where $P(N_{s}(\tau)=i|q;\theta)$ is the probability that a policy parameterized by $\theta$ will produce a trajectory $\tau$ with $N_{s}(\tau)=i$ given question $q$.
Given this, we will first prove the forward direction. Let $\theta^{*}$ be the parameters optimizing the objective and let $(q,y)$ be an arbitrary question-answer pair. Assume that $\pi_{\theta^{*}}$ must answer question $q$ parametrically. This firstly implies that $P(N_{s}(\tau)=0|q;\theta^{*})=1$ and, due to the optimality of $\theta^{*}$, that
$E_{\tau\sim\pi_{\theta^{*}}(\cdot|q,H)}[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)|N_{s}(\tau)=0]$ $=\max_{\theta}E_{\tau\sim\pi_{\theta}(\cdot|q,H)}[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)|N_{s}(\tau)=0]$
$=R^{*}_{0}(q,y).$
Furthermore, we note that the fact that $\theta^{*}$ must answer $q$ parametrically implies that
$R^{*}_{0}(q,y)=E_{\tau\sim\pi_{\theta^{*}}(\cdot|q,H)}[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)|N_{s}(\tau)=0]>R^{*}_{i}(q,y)$
for all $i>0$, where $R^{*}_{i}(q,y)$ is the maximum attainable expected reward when using $i$ searches. Assume for the sake of contradiction this were not true and that there existed some $i$ such that
$R^{*}_{0}(q,y)=E_{\tau\sim\pi_{\theta^{*}}(\cdot|q,H)}[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)|N_{s}(\tau)=0]\leq R^{*}_{i}(q,y).$
Let $\hat{\theta}$ be a policy that sets $P(N_{s}(\tau)=i|q;\hat{\theta})=1$ and displays the same behavior as $\theta^{*}$ all other questions. In the first case, $R^{*}_{0}(q,y)=R^{*}_{i}(q,y)$. Here, $\hat{\theta}$ would have the same expected reward as $\theta^{*}$ and be another optimal policy. However, this contradicts our assumption that the optimal policy must answer $q$ parametrically. In the other case, $R^{*}_{0}(q,y)<R^{*}_{i}(q,y)$. In this instance, $\hat{\theta}$ achieves a greater expected reward than $\theta^{*}$, which contradicts the assumption that $\theta^{*}$ is optimal. Hence, $R^{*}_{0}(q,y)>R^{*}_{i}(q,y)$ for all $i>0$, which proves the forward direction.
We will next prove the backwards direction. Let $\theta^{*}$ be the parameters optimizing the objective and assume that $R^{*}_{0}(q,y)>R^{*}_{i}(q,y)$. We claim that $\theta^{*}$ must answer $q$ parametrically. Let $C$ be an arbitrary set of non-negative integers containing at least one positive integer $i>0$ and assume for the sake of contradiction that $\theta^{*}$ set $P(N_{s}(\tau)=i|q;\theta^{*})>0$ for all $i\in C$ and $\sum_{i\in C}P(N_{s}(\tau)=i|q;\theta^{*})=1$. Then, it firstly follows that:
$E_{\tau\sim\pi_{\theta^{*}}(\cdot|q,H)}[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)]$ $=\sum_{i\in C}P(N_{s}(\tau)=i|q;\theta^{*})E_{\tau\sim\pi_{\theta^{*}}(\cdot|q,H)}[r_{acc}(y,\tau)\cdot r_{help}(q,\tau)|N_{s}(\tau)=i]$
$=\sum_{i\in C}P(N_{s}(\tau)=i|q;\theta^{*})R^{*}_{i}(q,y),$
due to our assumption that $\theta^{*}$ is optimal. However, due to our assumption that $R^{*}_{0}(q,y)>R^{*}_{i}(q,y)$ for all $i>0$, we also have that
$\sum_{i\in C}P(N_{s}(\tau)=i|q;\theta^{*})R^{*}_{i}(q,y)<\sum_{i\in C}P(N_{s}(\tau)=i|q;\theta^{*})R^{*}_{0}(q,y)=R^{*}_{0}(q,y).$
Given this, we can construct an alternative policy $\hat{\theta}$ that sets $P(N_{s}(\tau)=0|q;\hat{\theta})=1$ and behaves the same as $\theta^{*}$ on all other questions. This policy $\hat{\theta}$ will have a larger expected reward than $\theta^{*}$, which contradicts our assumption that $\theta^{*}$ is optimal. Hence, $\theta^{*}$ cannot set $P(N_{s}(\tau)=i|q;\theta^{*})>0$ for any $i>0$ and must answer $q$ parametrically. This proves the reverse direction. As a result, our claim holds.
Lemma 2:
Let $(q,y)$ be an arbitrary question-answer pair. Then, the maximum achievable expected accuracy when answering $q$ parametrically is equal to $R^{*}_{0}(q,y)$.
Proof of Lemma 2:
Consider an arbitrary trajectory $\tau$ such that $N_{s}(\tau)=0$. We claim that the accuracy of $\tau$ is equal to the reward of $\tau$. In the first possibility, $r_{acc}(y,\tau)=1$. Then, that the reward $r_{acc}(y,\tau)\cdot r_{help}(q,\tau)=1$ follows from Definition 2. In the second possibility, $r_{acc}(y,\tau)=0$. Then, the reward
$r_{acc}(y,\tau)\cdot r_{help}(q,\tau)=0\cdot r_{help}(q,\tau)=0.$
Hence, the accuracy and reward of $\tau$ are equal when $N_{s}(\tau)=0$. This further implies that the maximum achievable expected accuracy is equal to the maximum achievable expected reward when answer $q$ parametrically. This latter quantity, by Definition 3, is $R^{*}_{0}(q,y)$. The lemma is therefore correct.
Theorem:
For a given question-answer pair $(q,y)$, the optimal policy $\pi_{\theta^{*}}$ must answer the question $q$ parametrically (without any searches) if and only if the maximum achievable expected accuracy when answering $q$ parametrically is greater than or equal to the maximum achievable expected reward when performing searches.
Proof of Theorem:
Consider an arbitrary question-answer pair $(q,y)$ and let $\theta^{*}$ be the optimal policy. Then, Lemma 1 implies that $\theta^{*}$ must answer the question $q$ parametrically if and only if $R^{*}_{0}(q,y)>\max_{i>0}R^{*}_{i}(q,y)$. Furthermore, by Lemma 2, we know that the maximum achievable expected accuracy when answering $q$ is equal to $R^{*}_{0}(q,y)$. These two lemmas together imply that the optimal policy $\pi_{\theta^{*}}$ must answer the question $q$ parametrically if and only if the maximum achievable expected accuracy when answering $q$ parametrically is greater than or equal to the maximum achievable expected reward when performing searches. The theorem holds.
Appendix B Search Tool Use
@@ B.2 Inference Algorithm
0: Question $q$, language model $\pi_{\theta}$, retriever $H$
0: Maximum search budget $L$
0: Trajectory $\tau$
@@ B.3 Warm-Start
Evaluation of trajectories is done with an LLM judge, in this case Qwen2.5-72B-Instruct (Qwen et al., 2025). We follow the same procedure we use to evaluate abstention model outputs, described in Section C.1. If a trajectory is deemed correct, we swap its generated answer with the ground-truth answer for the target dataset to align answers with the dataset format, as we use exact match as the reward.
@@
Algorithm 2 Warm-Start Trajectory Construction
0: Datapoint $(q,a^{*})$, generator $G$, retriever $H$, maximum searches $l_{max}$, num samples NN
0: Datapoint $(q,\tau)$ for SFT
 Sample random number of searches $l\sim\{0,\ldots,l_{max}\}$
for $i=1\to N$ do
  Initialize current trajectory $\tau_{i}\leftarrow\emptyset$
  for action in seq do
   Append action start tag $\tau_{i}\leftarrow\tau_{i}~+<\text{action}>$
   Generate action $a\sim G(\cdot|q,\tau_{i})$ until <</action>>
   Append action $a$ to trajectory $\tau_{i}\leftarrow\tau_{i}+a$
   if action == search then
    Retrieve top-$k$ documents $o\sim H(a)$
    Append $o$ to trajectory $\tau_{i}\leftarrow\tau_{i}+o$
   end if
  end for
end for
return $\tau$
@@
B.4 Datasets
We run training experiments on three knowledge-intensive datasets – the single-hop dataset Natural Questions (Kwiatkowski et al., 2019), and multi-hop datasets HotPotQA (Yang et al., 2018) and 2WikiMultiHopQA (Ho et al., 2020). Table 7 displays the distribution of question answerability for each dataset. Specifically, each column represents the proportion of test set questions where Qwen2.5-3B-Base achieves an average accuracy of $i$, estimated using the procedure outlined in Appendix C.1. As seen in the table, the distributions are very similar for Natural Questions and HotPotQA, with $44.72\%$ and $50.11\%$ of questions having an average accuracy of 0 and $11.26\%$ and $7.92\%$ of questions having an average accuracy of $1$. 2WikiMultiHopQA is more challenging than both, with $77.26\%$ of questions having an average accuracy of 0 and only $0.85\%$ of questions having average accuracy of $1$.
We additionally use the single-hop TriviaQA dataset as part of our out-of-distribution evaluations. For Natural Questions, we use the official splits for training, validation and test. For HotPotQA, 2WikiMultiHopQA and TriviaQA, the official test splits do not contain answers. As a result, we use their official development/validation sets for the purpose of test and construct our own validation sets by sub-sampling from the training set with a 90/10 split. Additionally, as noted in the main text, we filter out the “comparison” and “bridge-comparison” questions from 2WikiMultiHopQA, as these questions are each binary choice questions with heavily skewed answer distributions, causing models to reward hack in practice.
Table 7: Proportion of test set questions where Qwen2.5-3B-Base achieves an average accuracy of $i$ for each $i\in\{0,0.1,\ldots,0.9,1\}$ for Natural Questions, HotPotQA and 2WikiMultiHopQA. We find that Natural Questions and HotPotQA have similar distributions of average accuracy, with 2WikiMultiHopQA being harder than both
Method Percentage of Questions with an Average Accuracy of $i$
0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
Natural Question 44.72 9.62 6.42 5.06 4.04 3.07 3.65 4.04 3.73 4.40 11.26
HotPotQA 50.11 9.41 5.91 4.75 3.86 3.94 3.54 3.47 3.71 3.37 7.92
2WikiMultiHopQA 77.26 8.49 3.77 2.46 1.97 1.33 1.22 0.96 0.85 0.83 0.85
Input Prompt:
Answer the given question. You should first have a reasoning process in mind and then provides the answer. Show your reasoning in <<think>> <</think>> tags and return the final answer in <<answer>> <</answer>> tags, for example <<answer>> Beijing <</answer>>. Question: <<question>>
@@ B.4 Datasets
Input Prompt:
Answer the given question. You must conduct reasoning between <<think>> and <</think>> every time you get new information. After reasoning, if you find you lack some knowledge, you can call a search engine by <<search>> query <</search>> and it will return the top searched results between <<document>> and <</document>>. You need to make every search call count and gain helpful results. If you find no further external knowledge is needed, you can directly provide the answer inside <<answer>> and <</answer>>, without detailed illustrations. For example, <<answer>> Beijing <</answer>>. Question: <<question>>
@@
Input Prompt:
Answer the given question. You must conduct reasoning between <<think>> and <</think>> every time you get new information. After reasoning, if you find you lack some knowledge, you can ask a question to a search engine by <<search>> query <</search>> and it will return the top searched results between <<document>> and <</document>>. A search query should be an atomic question asking about one, single piece of information.
Example 1: Question: “Who was born first, Clint Eastwood or Harrison Ford?” Valid Queries: “<<search>>Clint Eastwood birth date<</search>>” and “<<search>>Harrison Ford birth date<</search>>”. The query “<<search>>Clint Eastwood and Harrison Ford birth date<</search>>” is invalid. The query “
<<search>>
Clint Eastwood birth date
Harrison Ford birth date
<</search>>
is also invalid. Do not pack in multiple questions into one query. Each query should be completely independent.
Example 2: Question: “Which is a genus of palms, Zinnia or Butia?” Valid Queries: “<<search>>Zinnia genus classification<</search>>” and “<<search>>Butia genus classification<</search>>”.
Example 3: Question: “When did the country where Piltene is located become part of the USSR?” Initial Query: “<<search>>Piltene location<</search>>
In each of these examples, you should conduct a search only if you lack the relevant information. Remember, you should decompose questions in your search queries and conduct searches for each atomic question separately. You need to make every search call count and gain helpful results. If you find no further external knowledge is needed, you can directly provide the answer inside <<answer>> and <</answer>>, without detailed illustrations. For example, <<answer>> Beijing <</answer>>.
Question: <<question>>
@@
Appendix C Abstention Experiment Details
@@
C.1 Question Accuracy Estimation
We follow the pipeline used by  Yang et al. (2024) to estimate the average accuracies. For a given question $q$, we sample $10$ responses $\{\hat{y_{i}}\}^{10}_{i=1}$ from the untrained model. As all of our experiments are conducted with base models, we perform few-shot prompting. Specifically, for each dataset, we collect correct responses sampled from DeepSeek-V3.1 to $5$ questions sampled from the training set and use these as our few-shot examples. For this component, we perform inference with DeepSeek-V3.1 using a temperature of 11 and top-$p$ of $0.8$. We likewise perform sampling with Qwen2.5-3B with a temperature of $1$, top-$p$ of $0.8$ and top-$k$ of $50$ to ensure that the base model samples strong outputs and gives a good estimate of knowledge boundaries.
To assess the correctness of a given answer $\hat{y}_{i}$, we first extract a shortform response and then evaluate the accuracy of this extracted response with an LLM judge. We use DeepSeek-V3.1 in both cases using the few-shot prompts of Yang et al. (2024) (shown in Figures 7 and 8), using greedy decoding for replicability.
C.2 Trained Abstention Model Details
@@
C.3 Few-Shot Abstention Prompting Details
@@
Input Prompt:
Answer the given question. If you are not confident that your answer will be correct, you should abstain from answering by using the phrase ’I am afraid I cannot help you as I do not know the answer to this question.’ Question: <<question>>
Figure 5: The input prompt used during ternary abstention experiments. The final <<question>> is replaced by the input question.
C.4 Ternary Abstention Baseline Details
The Ternary baseline is trained with GRPO using a ternary reward that rewards correct answers with +1+1, abstentions with 0 and incorrect answers with 1-1. We assess correctness using exact match and assess abstention heuristically by checking whether a string from the list [i am afraid, i’m afraid, i cannot help, i can’t help, do not know, don’t know, abstain, abstention] appears as a substring in the final answer. Training uses the prompt shown in Figure 5. Similar to our oracle helper setting, we find that training with this ternary reward leads to models always abstaining within $25$ steps. This can be seen in Figure 6.
Figure 6: Abstention rate at different training steps when trained with a ternary reward for abstention. Models converge to always abstaining within 25 steps for all datasets.
C.5 Evaluating Abstention Models